OpenStudy (anonymous):

What is the product of m^2+2m-8/6p * 12p/4 m+m^2

OpenStudy (anonymous):

Assuming "m" is a variable Use "m^2" as a unit instead

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

if this is your problem: \(\large \frac{m^2+2m-8}{6p} \cdot \frac{12p}{4m+m^2} \) factor the numerator on that first fraction and the denominator of that last fraction then cancel out common factors...

OpenStudy (anonymous):

\(\huge \frac{m^2+2m-8}{6p} \cdot \frac{12p}{4m+m^2}=\frac{(m+4)(m-2)}{6p} \cdot \frac{12p}{m(m+4)} \)